OpenStudy (anonymous):

How many grams of sodium chloride can react with 18.3 liters of fluorine gas at 1.2 atmospheres and 299 Kelvin? F2(g) + 2NaCl(s)→CL2(g) + 2 NaFs The given information (volume, pressure and temperature) about the fluorine gas can be used to solve for moles of fluorine, which can then be used to start the stoichimetry problem. n(F2) = (1.2 atm)(18.3L)/(0.0821)(299 K) In this problem where is the 0.0821 coming from?

OpenStudy (jfraser):

0.0821 is the Ideal Gas Constant, R.\[P*V = n*R*T\]so \[R = \frac{P*V}{n*T}\] so R has units of \[R = 0.0821 (\frac{L*atm}{mol*K})\]